OpenStudy (anonymous):

does this series converge or diverge? n^(1/2)/(n+1)^3

OpenStudy (anonymous):

if it is \[\sum \frac{\sqrt{n}}{(n+1)^3}\] then it converges for sure

OpenStudy (anonymous):

I think you can compare with sqrt(n)/n^3, can't you?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\sqrt{n}/(n+1)^3\] explain!

OpenStudy (anonymous):

if you wnant a precise reason, do what bmp said if you want an easy quick eyeball answer the denominator has degree 3 and the numerator has degree \(\frac{1}{2}\) and \(3-\frac{1}{2}>1\) so it converges